OpenStudy (anonymous):

3√45x^3 + 2√12 + √27 - 3√20x^3

jhonyy9 (jhonyy9):

\[3\sqrt(3^{2}*5)x ^{2}*x + 2\sqrt(2^{2}*3 +\sqrt(3^{2}*3) -3\sqrt(2^{2}*5*x ^{2}*x\]

jhonyy9 (jhonyy9):

\[3*3\sqrt5x ^{3} +2*2\sqrt3 +3\sqrt3 -3*2\sqrt5x ^{3}\]

jhonyy9 (jhonyy9):

\[9\sqrt5x ^{3} +4\sqrt3 +3\sqrt3 -6\sqrt5x ^{2}\]

jhonyy9 (jhonyy9):

\[3\sqrt5x ^{3} +7\sqrt3\]